### v1

# Tally: Non-Intrusive Performance Isolation for Concurrent Deep Learning Workloads

Abstract.
GPU underutilization is a significant concern in many production deep learning clusters, leading to prolonged job queues and increased operational expenses. A promising solution to this inefficiency is GPU sharing, which improves resource utilization by allowing multiple workloads to execute concurrently on a single GPU. However, the practical deployment of GPU sharing in production settings faces critical obstacles due to the limitations of existing mechanisms, such as high integration costs, inadequate performance isolation, and limited application compatibility. To address these issues, we introduce Tally, a non-intrusive GPU sharing mechanism that provides robust performance isolation and comprehensive workload compatibility. Tally operates as a virtualization layer between applications and GPUs, transparently orchestrating the device execution of concurrent workloads. The key to Tally’s robust performance isolation capability lies in its fine-grained thread-block level GPU kernel scheduling strategy, which allows the system to effectively mitigate interference caused by workload co-execution. Our evaluation, conducted on a diverse set of workload combinations, demonstrates that Tally on average incurs a mere $7.2\%$ overhead on the $99^{th}$-percentile latency of high-priority inference tasks when executed concurrently with best-effort training workloads compared to $188.9\%$ overhead exhibited by the state-of-the-art GPU sharing systems like TGS, while achieving over $80\%$ of TGS’s system throughput.
Part of this work was done while the author was a student at the University of Toronto.
1. INTRODUCTION
Deep learning (DL) has demonstrated astonishing performance across a wide range of applications, including image recognition (krizhevsky2012, ; kaiming2016, ; szegedy2015, ), machine translation (devlin-etal-2019-bert, ; bahdanau2015neural, ; sutskever2014, ; lewis-etal-2020-bart, ), recommendation systems (Zhang2020, ; alashkar2017examples, ), and autonomous driving  (Cordts2016, ; yurtsever2020survey, ). Recent advances in generative AI (rombach2022high, ; brown2020language, ), exemplified by the development of large language models (LLMs) like ChatGPT (openai2024chatgpt, ), have further elevated the influence of DL to unprecedented levels. To facilitate the research and deployment of DL-based applications, many organizations (shukla2022singularity, ; Weng2022mlaas, ; meta2024, ) are investing in building computing infrastructures with large fleets of GPU devices. For example, Meta is building two data center-scale clusters for AI, each equipped with hundreds of thousands of NVIDIA H100 GPUs with an estimated cost as high as half a billion dollars (meta2024, ; Eadline_2023, ). As the demand for computational resources continues to escalate, efficient utilization of GPUs becomes paramount to mitigate costs and ensure scalability for AI applications.
Nevertheless, studies have noted significant GPU underutilization in large-scale DL clusters. For example, a recent study by Alibaba (Weng2022mlaas, ) revealed that the median GPU utilization in their production DL cluster can go as low as $4.2\%$. This inefficiency results from the need for production clusters to accommodate a diverse range of workloads, many of which are unable to fully leverage the capacity of the GPUs. For instance, training processes can potentially underutilize GPUs due to bottlenecks in CPU execution such as stalls caused by data fetching and preprocessing (Audibert2023, ). The utilization of online inference tasks, on the other hand, is intrinsically tied to real-time request rates, which are subject to substantial fluctuations causing unpredictable GPU idleness (li2023-alpaserver, ). For instance, Microsoft’s production trace (maf2, ) has revealed significant burstiness in user requests, with frequent spikes in demand of up to 50×50$\times$50 × the average. This also makes usage prediction and resource allocation challenging for inference workloads, resulting in resource over-provisioning to ensure high quality of service (MArk2019, ; li2023-alpaserver, ), which exacerbates the underutilization problem.
To enhance the efficiency of DL clusters, recent studies (xiao2020, ; salus2020, ; lucid, ; narayanan2020heterogeneity, ) suggest GPU sharing as a promising solution. This approach aims to enhance hardware utilization by allowing multiple low-GPU tasks to share a single GPU’s resources. The same Alibaba study conducted a simulation on their production traces and demonstrated that an effective GPU sharing mechanism could potentially reduce the amount of GPU resources required by the cluster by $50\%$ on average and up to $73\%$ during peak hours. These findings underscore the potential of GPU sharing in mitigating the GPU underutilization problem and significantly improving the overall efficiency of large-scale DL clusters.
Despite these promising results, GPU sharing has not yet achieved widespread adoption in production. One major obstacle is the lack of performance isolation mechanisms in most current GPU sharing systems. Many workloads within DL clusters are governed by service-level agreements (SLAs), which impose stringent performance requirements, usually on the order of milliseconds (clipper2017, ). These workloads are commonly referred to as high-priority or latency-critical tasks. By contrast, workloads that can tolerate extended processing times are known as low-priority or best-effort tasks. While it seems logical to pack best-efforts tasks along with high-priority tasks to utilize the GPU idle time caused by the latter, this approach can cause significant interference to the performance of the high-priority workloads, potentially violating their SLAs. Our experiments on state-of-the-art GPU sharing mechanisms reveal that such performance interference can cause up to a 20×20$\times$20 × slowdown in tail latency (see Section 5), making them impractical for use in practice.
Another hindrance to the broader adoption of GPU sharing lies in the intrusive nature of many current sharing solutions  (salus2020, ; xiao2020, ; case2022, ; effisha2017, ; slate2019, ; tacker2022, ). Using such solutions require extensive modifications to user-level code or the underlying DL frameworks (e.g., PyTorch  (paszke2019pytorch, ), TensorFlow (tensorflow2016, )). These solutions not only greatly limit the flexibility of users in developing their workloads but also impose a substantial burden on the users and the cluster providers for workload adaptation and framework upgrades, leading to significant integration and maintenance costs (lucid, ). Moreover, many of these solutions require the workloads to adhere to specific characteristics, such as ensuring GPU kernels used by the workload are always idempotent (reef2022, ). These constraints often only apply narrowly in production clusters that host a diverse range of workloads with varying characteristics.
In light of these observations, we argue that for practical deployment in production DL clusters, a GPU sharing system must meet the following three criteria. (1) Non-intrusive to seamlessly integrate into existing clusters and DL workloads, (2) provide performance isolation guarantees to ensure that the service-level requirements of high-priority tasks are maintained effectively during shared execution, and (3) be generalizable across a diverse range of DL applications. Building on these objectives, we propose Tally, a novel non-intrusive GPU sharing system for large-scale production DL clusters that ensures performance isolation of latency-critical high-priority tasks when co-located with best effort tasks, while maximizing hardware efficiency and overall system throughput. We design Tally as a transparent virtualization layer between the application-level and the GPU. It intercepts GPU kernel launch API calls and employs a task-agnostic scheduling algorithm to prioritize high-priority kernels, while opportunistically scheduling best-effort kernels during GPU idle periods.
To efficiently co-locate high-priority and best-effort tasks on a GPU with performance isolation guarantees, it is essential to schedule best-effort kernels in a controlled manner when the GPU is idle and promptly switch to high-priority kernels as required. However, modern GPUs lack kernel scheduling primitives to terminate a running kernel and only allow context switching between kernels at the granularity of entire kernels (reef2022, ; GILMAN2021102234, ). We observe that this limitation introduces significant overhead on high-priority tasks, especially since many best-effort DL workloads involve long-running kernels. To overcome this challenge and enable finer-grained kernel scheduling, Tally uses two software-level kernel scheduling primitives: slicing and preemption. Slicing divides large kernels into smaller segments, while preemption allows for the interruption of active kernels. Our key insight is that these two scheduling primitives can be implemented non-intrusively and in a task-agnostic manner through a series of GPU kernel transformation passes, leveraging the GPU programming model which organizes kernels as a collection of independent thread blocks.
While slicing and preemption are effective mechanisms to manage co-execution interference, each has its trade-offs. Kernel slicing is a lightweight transformation but can incur higher overhead due to multiple kernel launches. In contrast, kernel preemption requires only a single kernel launch but involves a more complex transformation with potential synchronization overhead. Due to these differences, the optimal choice of scheduling primitives and configuration parameters (e.g., the degree of slicing) varies across different kernels. Towards this end, Tally uses a profile-guided resource provisioning strategy in its scheduling algorithm that transparently profiles the performance of best-effort kernels on-the-fly across various scheduling configurations to select the best option to meet the latency requirements of the high-priority task. This approach ensures that interference from best-effort tasks remains within acceptable limits while striving for optimal overall system throughput.
To evaluate Tally’s ability to provide effective performance isolation, we prepare a comprehensive benchmark suite consisting of six training and six inference workloads that cover a wide range of popular DL models and frameworks (details shown in table 2). We evaluate the co-located performance of latency-critical inference tasks and best-effort training tasks by simulating the production traffic patterns for the inference tasks using the publicly available Microsoft Azure Function Trace 2021 (MAF2) dataset (maf2, ). Across all the combinations, Tally incurs a mere $7.2\%$ overhead on average in the $99^{th}$-percentile latency of inference tasks, compared to $252.3\%$, $345.0\%$, $195.5\%$, and $188.9\%$ exhibited by state-of-the-art non-intrusive GPU sharing solutions, Time-Slicing, MPS, MPS-Priority, and TGS respectively. At the same time, Tally is able to achieve $80\%$ of the system throughput attained by TGS, demonstrating the effectiveness of Tally in maximizing the GPU efficiency while providing strong performance isolation guarantees.
In summary, we make the following contributions.
  • We propose Tally, a non-intrusive GPU sharing solution with robust performance isolation guarantees.
  • We introduce a novel approach to achieve fine-grained GPU kernel scheduling in a non-intrusive and task-agnostic manner using kernel slicing and preemption primitives. We further propose a priority-aware kernel scheduler that dynamically selects the appropriate scheduling primitives to optimize GPU throughput while preserving performance isolation guarantees.
  • We evaluate Tally on a diverse range of inference and training workloads and show that Tally can achieve system throughput on-par with state-of-the-art GPU sharing solutions while providing strong performance isolation guarantees.
2. BACKGROUND
Modern large-scale DL clusters are typically shared among many users within an organization and are designed to support a diverse range of workloads (xiao2020, ; lucid, ), encompassing training, fine-tuning, and inference tasks for a variety of model architectures, such as convolutional neural networks (CNNs), transformer-based large language models (LLMs), graph neural networks (GNNs), and reinforcement learning (RL) models. These tasks are often implemented using various mainstream DL frameworks such as PyTorch, TensorFlow, and JAX, as well as proprietary softwares (Weng2022mlaas, ). Given the diversity and compute-intensive nature of DL workloads, general-purpose GPUs have become the dominant accelerator in DL clusters due to their powerful programming model and immense parallel computing capabilities.
GPU architecture and programming model. While there are variations among vendors (e.g, NVIDIA and AMD), modern GPUs used in DL clusters fundamentally follow similar architecture and programming model. GPUs are equipped with thousands of parallel processing units called Streaming Multiprocessors (SMs). Kernels, the computational units executed on GPUs, are programmed using specialized languages like CUDA (cuda2024, ). GPU kernels follow a hierarchical thread structure consisting of threads, warps, thread blocks, and grids. Threads within a kernel identify themselves using built-in variables such as threadIdx and blockIdx to determine their task assignments. Warps are groups of threads that execute in SIMT (Single Instruction Multiple Thread) fashion. All the warps within a thread-block are scheduled to a single SM and can cooperate using fast on-chip shared memory. Additionally, threads within a block can synchronize their execution to coordinate memory accesses using synchronization primitives. Finally, thread blocks within a grid can be scheduled and executed independently in any order, either sequentially or in parallel (cuda2024, ). This independent scheduling capability allows thread blocks to scale effectively across multiple SMs.
GPU sharing systems. The diverse nature of DL workloads, coupled with their inherent inefficiencies, has raised significant concerns about GPU underutilization in many production DL clusters (hu2021characterization, ; xiao2020, ). To mitigate this problem, researchers have proposed numerous cluster-level optimizations (xiao2018, ; xiao2020, ; wang2021horizontally, ; salus2020, ), among which GPU sharing is a promising approach to improve utilization by consolidating multiple low-utilization jobs onto the same GPU. From an implementation perspective, current GPU sharing solutions can be broadly categorized into two levels: application and runtime. Application-level solutions propose a new programming interface for writing efficient DL workloads that can be executed concurrently. For example, Salus (salus2020, ) exposes a set of scheduling primitives for defining the iteration boundaries of a training job, enabling the scheduler to manage different training workloads at the iteration-level granularity. Similarly, HFTA (wang2021horizontally, ) provides APIs that allow developers to write multiple hyper-parameter tuning jobs as a single application, which can then be efficiently executed on a single GPU.
Unlike application-level solutions that necessitate modifications to user-level applications, runtime-level GPU sharing solutions operate transparently by integrating directly into either DL frameworks (xiao2020, ) (e.g., PyTorch (paszke2019pytorch, ) or TensorFlow (tensorflow2016, )), container orchestration systems (tgs2023, ) (e.g., Kubernetes (kubernetes, )), or the GPU driver (reef2022, ). These systems typically perform resource sharing at a more fine-grained level. For instance, TGS (tgs2023, ) is designed to be deployed on GPU containers in Kubernetes and schedules workloads at the GPU kernel level, while REEF (reef2022, ) extends GPU drivers to offer more fine-grained thread-level scheduling capabilities. The transparent nature of many these solutions allows for adoption across a wider range of applications with minimal effort. However, designing these solutions is often challenging due to the lack of application-level context.
3. GPU SHARING IN THE WILD
Despite the wide range of solutions, current GPU sharing systems are inadequate for practical deployment in DL clusters due to one or more of the following key limitations.
High integration and maintenance cost. Many existing GPU sharing systems are fundamentally intrusive in nature, necessitating extensive modifications to user code to be used in practice. This is particularly evident in application-level systems, where users must adapt their workloads to the specific programming interfaces proposed by these solutions. Such intrusive approaches not only burden the users, but also limit their flexibility in application development (lucid, ). Similarly, runtime-level GPU sharing solutions that are designed as extensions of DL frameworks impose a significant burden on cluster managers to implement and maintain them across diverse frameworks used by the users. Consequently, deploying these solutions in practice often involves substantial integration and maintenance costs, especially given the rapidly evolving landscape of the modern DL ecosystem.
On the other hand, runtime-level GPU sharing solutions that seamlessly integrate with the container orchestration system or GPU driver, although may seem non-intrusive, they often impose implicit requirements on user code to gather additional application context needed for making effective scheduling decisions. This include requirements such as enabling the ability to temporarily launch applications in a specialized profiling environment before execution (orion2024, ), or mandating applications to implement checkpoint and restore mechanisms (gavel2020, ; xiao2018, ). Enforcing such programming practices in large-scale multi-tenant clusters is often impractical (Weng2022mlaas, ).
Lack of performance isolation guarantees. In production DL clusters, many inference tasks are governed by service-level agreements (SLAs) which impose strict latency requirements on their response times, typically of the order of a few milliseconds to seconds (clipper2017, ). However, current GPU sharing solutions often lack support for performance isolation, resulting in unacceptable performance degradation for latency-critical inference tasks when co-located with other tasks (GILMAN2021102234, ). For instance, NVIDIA’s native runtime-level GPU sharing solution, MPS (nvidia-mps, ), is designed to eagerly schedule as many GPU kernels from co-located tasks as possible. Although this design maximizes the GPU utilization, it can cause significant queuing delays for executing kernels of high-priority inference tasks. In such scenarios, our experiments (see Section 5) indicate that, the tail latency of a co-located inference task can go as high as 20×20$\times$20 × its standalone performance.
Effectively sharing GPU resources while ensuring performance isolation for latency-critical tasks requires the ability to prioritize one task over the other and rapidly switch between them based on their resource demands. The speed of context-switching between tasks, which we denote as turnaround latency, is dictated by the scheduling granularity used in GPU sharing systems. As described in Section 2, the scheduling granularity used by prior works primarily fall into four categories, from coarsest to finest: iteration-level (e.g. Salus (salus2020, )), kernel-level (e.g., TGS (tgs2023, )), block-level (e.g. EffiSha (effisha2017, )), and thread-level (e.g. REEF (reef2022, )). To illustrate the effect of scheduling granularity on the the tail latency, we present a representative empirical study in Table 1 where a latency-critical BERT (devlin-etal-2019-bert, ) inference task is co-executed with a Whisper (radford2023robust, ) model training workload on an NVIDIA A100 GPU. In this experiment, we empirically estimate the turnaround latency for different scheduling granularities on the Whisper model and compare it to the response time of BERT inference. Since BERT inference takes only $3.93$ ms, providing strong performance isolation of such tasks require sub-millisecond-level turnaround latency. However, both iteration-level and kernel-level scheduling can only achieve turnaround latencies of $3$ s and $10$ ms, respectively. As a result, scheduling the Whisper training task at these granularities can introduce 2.5760×2.5-760$\times$2.5 - 760 × higher overhead on BERT’s latency. In contrast, block-level and thread-level scheduling achieve microsecond-scale turnaround latencies of $304$ µs and $38$ µs respectively resulting in only mere $1-7\%$ overhead.
From these results, we conclude that to achieve practical performance isolation guarantees, GPU sharing solutions must schedule resources at either the block-level or thread-level granularity. Unfortunately, existing solutions either do not meet these criteria (e.g., AntMan, TGS) or prove ineffective in diverse production DL clusters due to their intrusive nature (e.g., EffiSha, Orion) or lack of generalizability (e.g. REEF). This brings us to the third limitation.
Table 1. Comparison of BERT inference latency against turnaround latency of different scheduling granularity for Whisper training on NVIDIA A100 GPU.
Inference time (BERT) Turnaround latency (Whisper)
Iteration Kernel Block Thread
$3.93ms$ $\sim 3s$ $\sim 10ms$ $\sim 304\mu s$ $\sim 38\mu s$
Reliance on narrowly applicable workload characteristics. Many state-of-the-art GPU sharing systems rely on specific workload characteristics in their scheduling strategies. For instance, AntMan and TGS can provide runtime-level GPU sharing support but are limited to training workloads with predictable GPU kernel execution patterns. REEF, on the other hand, offers strong performance isolation for inference workloads, however, needs the GPU kernels to be idempotent to support thread-level kernel scheduling strategy, i.e., the kernel must always produce the same output for the same input, regardless of how many times it is executed. While these systems demonstrate benefits for certain class of applications, their underlying assumptions are often not applicable to large-scale multi-tenant DL clusters that must support a diverse range of workloads with varying characteristics. Prior studies (Weng2022mlaas, ) have revealed that DL clusters are often agnostic to the specific semantics of running workloads and extracting application-level context is largely infeasible. As a result, GPU sharing solutions that depend on specific task characteristics are not viable in many DL clusters.
In summary, we argue that a GPU sharing solution for large-scale DL clusters must address all three of the aforementioned limitations.
4. TALLY: SYSTEM DESIGN
To facilitate the practical deployment of GPU sharing in production clusters, we propose Tally, a novel non-intrusive GPU sharing system that offers robust performance isolation for latency-critical tasks during shared execution with throughput-oriented best-effort tasks on a diverse range of DL workloads. Tally is designed as a virtualization layer operating transparently between applications and GPUs, intercepting device API calls and employs a task-agnostic scheduling algorithm to prioritize execution from high-priority tasks while opportunistically scheduling kernels from best-effort tasks during GPU idle cycles.
Tally ensures performance isolation for high-priority tasks by scheduling kernels from best-effort tasks at the block-level granularity using two scheduling primitives: slicing and preemption. Slicing divides a large kernel into numerous smaller segments, enabling more controlled scheduling, whereas preemption allows for the interruption of active kernels. Importantly, we make two key observations that allows applying these two scheduling primitives in a task-agnostic and non-intrusive manner. First, unlike thread-level scheduling, thread blocks within a GPU are guaranteed to operate independently, making block-level scheduling universally applicable across GPU kernels. Moreover, while thread-level scheduling can provide stronger performance isolation, we find that block-level scheduling is sufficient for most practical scenarios. Second, the slicing and preemption scheduling primitives can be applied on GPU kernels through a series of transformation passes on the kernel device code (PTX in case of NVIDIA GPUs) obtained through the interception of device API calls. Leveraging these insights, we integrate the slicing and preemption primitives to Tally’s scheduling algorithm enabling microsecond-scale turnaround latency, allowing Tally to promptly switch execution between high-priority and best-effort tasks and maintain the millisecond-scale latency requirements in many latency-critical tasks.
Figure 1. Tally architecture
Figure 1 shows the high-level design of Tally, following a client-server architecture. On the application side, Tally’s client library intercepts device code and API calls initiated by client processes and forwards them to the Tally server, which is responsible for managing the actual device execution and scheduling. The objective of the Tally server is to facilitate priority-enforced concurrent execution of one high-priority task alongside multiple best-effort tasks. To achieve this, Tally server incorporates three key components: a (1) kernel transformer, which leverages intercepted device code to convert kernels into slicing and preemptive-style kernels, a (2) transparent profiler, which profiles the performance of the transformed kernels under various launch configurations at runtime and predicts their estimated turnaround latencies, and finally a (3) priority-aware scheduler, which integrates capabilities of the aforementioned modules and strategically schedules GPU execution of concurrent processes in a priority-aware manner. In the following sections, we provide details regarding each of these components.
Figure 2. Kernel Transformations: (a) Slicing and Preemption Transformation (b) Unified Synchronization Transformation
4.1. Kernel Transformations
We first illustrate how Tally supports the slicing and preemption primitives through a series of transformation passes on the kernel device code.
Slicing transformation. Slicing is a technique that facilitates fine-grained control over kernel execution by partitioning the thread blocks of a kernel into multiple sub-kernels. Figure 2(a) illustrates this concept, where a kernel originally composed of $16$ thread blocks is divided into $4$ sub-kernels, each containing $4$ blocks. As thread blocks within a kernel are guaranteed independent, theoretically, this partitioning should preserve the original kernel’s functional semantics. However, naively launching the original kernel $4$ times with $4$ thread blocks each will yield incorrect results, as GPU threads rely on built-in variables such as blockIdx to determine their task assignment. Specifically, the blockIdx variable informs a thread of the index of the block it resides in within the kernel execution, allowing it to properly compute its assigned task index. In the original kernel, threads are exposed to blockIdx values ranging from 00 to $15$ (as $16$ blocks in total are launched). Yet, when each sub-kernel is launched, the threads are exposed to blockIdx values ranging from 00 to $3$, causing all sub-kernels to execute the first $4$ blocks of the kernel, which will lead to erroneous results. To address this issue, we modify a kernel’s PTX code to take in an additional offset variable in its parameters and substitute occurrences of blockIdx with the sum of the offset and the original blockIdx. This approach enables threads within a sub-kernel to identify their corresponding subset of blocks. For instance, sub-kernel $2$ can reconstruct its corresponding block indices from $4$ to $7$ by adding a block offset of $4$ to the new blockIdx. With this mechanism, the collective work done by the sub-kernels can properly resemble the original kernel execution, while offering the scheduler flexible scheduling at the block-level granularity.
Preemption transformation. Tally introduces a transformation that makes kernels preemptible while preserving their functional semantics. The preemption transformation in Tally is inspired by the Persistent Thread Block (PTB) programming paradigm, which deviates from the traditional programming model where the number of thread blocks launched by a kernel is proportional to the amount of work it performs. Instead, the PTB model employs a flexible number of worker blocks to perform all the work. As an example in Figure 2(a), rather than launching $16$ thread blocks of the original kernel, the PTB model launches $2$ worker blocks to process the $16$ blocks in an iterative and dynamic manner. These workers continuously fetch and increment a global integer representing the next task index. Similar to how sub-kernels in slicing utilize block offsets to recompute their block indices from the original kernel execution, workers in PTB use the task index to reconstruct the corresponding block indices for the current iteration. For example, in the illustrated case where thread blocks are structured in a $1$-dimensional grid, a fetched task index of $8$ indicates that the worker should perform the execution of the thread block with blockIdx $8$. In the case of higher-dimensional grids, block indices can be computed from the integer task index by referring to the original kernel’s grid dimension.
A critical observation from the PTB model is that the progress of the kernel execution is effectively encapsulated in the global task index. Since the execution is now transformed into an iterative approach, kernels can potentially terminate early and later be resumed from the point at which they ended, using the task index. This can be achieved by introducing an additional boolean variable to the PTB kernels as a flag for signaling preemption. At the start of each iteration, workers check this flag to detect preemption signals, allowing block-level preemption within kernels. In Tally, we automatically convert kernels into this preemptive style by wrapping the original kernel function in the device code with an outer loop using control flow logic. For instance, all return statements are transformed into branch instructions that direct the execution flow to the start of the iteration, ensuring that each worker block continues execution until all tasks are processed or a preemption signal is received. Furthermore, a synchronization call is introduced at the end of each iteration to ensure that threads within a worker block process a single block at a time, preventing potential race conditions and maintaining correctness of computation.
However, since the preemption transformation replaces return statements with branching statements that synchronizes threads at the start of the iteration, it can potentially lead to threads within the same block synchronizing at different places. Such divergence in synchronization is an undefined behavior in the GPU programming model and, in practice, results in infinite kernel stalls, making such transformations unsafe. To address this issue, Tally introduces a prepositional kernel transformation pass, which we denote as the unified synchronization transformation. The algorithm of the transformation is depicted in Figure 2(b). As illustrated, this transformation modifies all synchronization and return instructions by redirecting them to a newly introduced unified synchronization point that verifies if all threads are in a return state. Threads from synchronization points loop back to their previous positions, while those returning are held until all threads align. This unified synchronization transformation eliminates divergences in thread block synchronization, and in turn, the infinite stall issue. Through this prepositional transformation stage, the preemptive transformation can be applied automatically and safely.
{minted}
[ linenos, fontsize=, xleftmargin=12pt, numbersep=5pt, ]python def launch_and_profile(kernel): candidates = get_candidate_configs(kernel) for cfg in candidates: est_turnaround = lookup_measurement(kernel, cfg) if est_turnaround is None: est_turnaround = profile(kernel, cfg) if has_launched(): return set_launch_config(kernel, candidates, bound=TUNEAROUND_LATENCY_BOUND) def scheduler(): while True: for client in sort_by_priority(clients): if client.is_high_priority(): kernel = client.fetch_next_kernel() if kernel is not None: preempt_best_effort_kernel() kernel.launch(Config::DEFAULT) if client.has_kernel_running(): break else: kernel = client.get_curr_ex_kernel() if kernel is None: kernel = client.fetch_next_kernel() cfg = lookup_launch_config(kernel) if cfg is not None: kernel.launch(cfg) else: launch_and_profile(kernel) elif kernel.running(): pass elif kernel.stopped(): kernel.resume()
Figure 3. Tally priority-aware scheduling algorithm.
4.2. Priority-aware Scheduling
We now demonstrate how Tally’s priority-aware scheduler leverages the aforementioned slicing and preemption primitives and the transparent profiler to enforce priority among the scheduling of concurrent workloads. The algorithm of the scheduler, as outlined in Figure 3, follows an opportunistic strategy: it prioritizes the immediate dispatch of high-priority kernels upon their arrival (lines $14$ to $20$) and schedules low-priority kernels to execute when the high-priority job becomes inactive (lines $21$ to $33$). Due to the reactive nature of this algorithm, incoming high-priority kernels may need to wait for the completion of currently executing low-priority kernels, potentially introducing delays to the execution of high-priority tasks. To mitigate such queuing delays, each best-effort kernel is either launched as multiple sub-kernels one at a time, or in its preemptive version, such that they can be preempted in the event of a high-priority kernel arriving during its execution.
The primary challenge of this strategy lies in determining the appropriate launch configuration for each best-effort kernel, which includes deciding whether to employ slicing or preemption and setting the corresponding parameters for each method. First, due to the inherent difference in their implementation, each of the two primitives exhibits distinct performance trade-offs. While slicing is a lightweight transformation, it can incur overhead due to multiple kernel launches. Conversely, the preemption kernel requires only a single launch, but the additional synchronization and control flow can introduce noticeable overhead to certain kernels. As a result, their relative performance across different kernels varies considerably, posing challenges for the scheduler in choosing between the two. Moreover, for each primitive, the scheduler need to determine the corresponding parameter, which includes the number of slices for slicing and the number of workers for preemption. These decisions are crucial as they significantly affect the performance behavior of the concurrent workloads. For instance, in the context of preemptive scheduling, allocating fewer workers can expedite the preemption process but can potentially slow down the execution of best-effort tasks, thereby diminishing the overall benefits of GPU sharing.
[…]
4.3. Non-intrusive Virtualization Layer
Finally, we present details on the implementation of Tally’s virtualization layer. On the application-side, Tally utilizes the LD_PRELOAD mechanism in Linux to intercept device API calls, redirecting them to a surrogate function that mirrors the API’s original signature. These intercepted API calls are sent to the server, which handles the actual device execution. The challenge in this process is enabling the server, which operates in a separate address space, to execute client kernel functions. The solution leverages a critical insight: GPU applications register device code with the GPU driver at the start of execution. By intercepting this process, the server gains access to the device code, including kernel binaries and device code, which we use at the server-side to apply kernel transformations and recompile to executable GPU kernels.
A notable concern with this interception and forwarding approach is the potential communication overhead, which could impact application performance. To mitigate this, we have implemented two critical optimizations. First, we utilize efficient shared memory channels for client-server communication, which significantly reduces the overhead by avoiding context switches during message passing. Second, we observe that the need to forward many frequent device API calls, such as cudaGetDevice in NVIDIA, can be avoided by maintaining local states of the execution context on the client side. This approach effectively minimizes the frequency of API call forwarding between the client and server, further reducing the communication overhead. These strategies collectively enable Tally to operate transparently with near-native performance across all the concurrent DL tasks.
5. EVALUATION
5.1. Methodology
Workloads. For Tally evaluation, we prepare a benchmark suite comprising six training and six inference workloads as listed in Table 2, covering a diverse range of widely used DL model architectures, such as CNNs (e.g., ResNet50 (kaiming2016, )), transformer-based LLMs (e.g., Llama-2 (touvron2023llama, )), and diffusion models (e.g., Stable Diffusion (rombach2022high, )). Training workloads are implemented using PyTorch (paszke2019pytorch, ), a broadly adopted DL framework in the research community. For inference workloads, we use high-performance engines such as ONNX Runtime (onnxruntime, ), TorchInductor (torchinductor2024, ), and Hidet (hidet2023, ). Inspired by prior DL serving benchmarks (li2023-alpaserver, ; reef2022, ; bhattacharjee2019barista, ), we simulate inference workloads using the Microsoft Azure Function Trace 2021 (MAF2) (maf2, ), a dataset of Azure serverless function invocations that has been repurposed for DL serving workloads. We extract the trace of the most frequently invoked function to adapt it to match specific traffic loads based on the query latency of the models. We define load as the percentage of time an inference service is actively serving requests.
Metrics. We use the $99^{th}$ percentile latency as the primary metric for assessing the performance of inference tasks. For both training and inference tasks, we measure their throughput, defined as the number of samples processed per unit of time. To evaluate the overall performance of concurrent task execution, we utilize system throughput, calculated by aggregating the normalized throughput of individual workloads.
  Training Model Dataset # Params Throughput
ResNet50 (kaiming2016, ) ImageNet (imagenet2009, ) 25.6M 1.0 it/s
PointNet (Qi_2017_CVPR, ) ShapeNet (chang2015shapenet, ) 3.5M 40.0 it/s
BERT (devlin-etal-2019-bert, ) SQuAD (rajpurkar-etal-2016-squad, ) 110M 1.8 it/s
GPT2-Large (radford2019language, ) Wikitext2 (merity2016pointer, ) 774M 3.3 it/s
PEGASUS (zhang2020pegasus, ) XSum (xsum-emnlp, ) 568M 2.9 it/s
Whisper-v3 (radford2023robust, ) LibriSpeech (librispeech2015, ) 1.5B 0.3 it/s
      Inference Model Engine # Params Latency
ResNet50 (kaiming2016, ) Hidet (hidet2023, ) 25.6M 1.37 ms
BERT (devlin-etal-2019-bert, ) ONNX RT (onnxruntime, ) 110M 3.93 ms
YOLOv6m (li2022yolov6, ) TorchInductor (torchinductor2024, ) 34.9M 17.5 ms
Llama-2-7b (touvron2023llama, ) ONNX RT (onnxruntime, ) 7B 1.9 s
Stable Diffusion (rombach2022high, ) TorchInductor (torchinductor2024, ) 983M 2.5 s
GPT-Neo (gpt-neo, ) TorchInductor (torchinductor2024, ) 2.7B 3.6 s
Table 2. Training and inference benchmarks
Figure 4. Analysis of 99th-percentile latency and system throughput across different high-priority inference and best-effort training workload combinations under various GPU sharing mechanisms.
Figure 5. (a) Latency and throughput for high-priority BERT and Llama2-7B inference co-located with BERT, GPT-2, and Whisper training under Tally and TGS across different traffic loads. (b) Time-series visualization of user traffic, tail latency, and throughput over time for BERT inference co-located with BERT training across different GPU sharing systems.
Baselines. We compare the performance of Tally against four state-of-the-art non-intrusive GPU sharing solutions: (i) Time-Slicing, (ii) MPS, (iii) MPS-Priority and (iv) TGS. All these solutions perform scheduling at the kernel-level granularity. More specifically, Time-Slicing and MPS are the proprietary concurrency solutions provided by NVIDIA GPUs. MPS-Priority refers to MPS with the client priority control feature enabled, which allows specifying priority levels among concurrent workloads. Lastly, TGS is a GPU sharing system specifically designed for DL workloads, which leverages adaptive rate control mechanisms to achieve performance isolation among concurrent workloads.
Experimental setup. All experiments are conducted on an AWS EC2 p4d.24xlarge instance, a production-grade GPU server widely deployed in DL clusters. This machine has $96$-core Intel Xeon Platinum 8275CL CPU, $1152$ GB of DRAM, and $8$ NVIDIA A100 GPUs, each with $40$ GB of memory. The server operates on Ubuntu 20.04 with CUDA 12.2, PyTorch 2.2.0, ONNX Runtime GPU 1.17.0, and Hidet 0.3.0.
5.2. End-to-end Results
We first analyze Tally’s end-to-end performance in efficiently sharing GPU resources among multiple workloads while guaranteeing performance isolation for high-priority inference tasks. To this end, we measure the $99^{th}$-percentile latencies of the inference tasks and the overall system throughput achieved by Tally when executed concurrently with a best-effort training workload. First, we compare the latency metrics against the state-of-the-art non-intrusive GPU sharing systems, namely Time-Slicing, MPS, MPS-Priority, and TGS, as well as against the ideal scenario of running the inference task in isolation. Figure 4 presents the latency and throughput measurements for all inference-training combinations from the benchmark suite, with inference tasks simulated using the MAF2 trace with a fixed $50\%$ load. The results demonstrate Time-Slicing, MPS, MPS-Priority, and TGS exhibit significantly high overhead compared to the ideal latency, with an average of $252.3\%$, $345.0\%$, $195.5\%$, and $188.9\%$, respectively. The suboptimal inference latencies exhibited by Time-Slicing and MPS can be attributed to their priority-agnostic design, which focuses solely on maximizing system utilization. MPS-Priority and TGS, on the other hand, exhibit improved latency by prioritizing kernel execution of high-priority tasks over best-effort tasks. However, their coarse-grained kernel-level scheduling strategy fails to effectively mitigate co-execution interference. In stark contrast, Tally consistently maintains tail latencies nearly identical to the ideal latencies of the high-priority workloads, with an average percentage increase of mere $7.2\%$. This stronger performance isolation capability can be attributed to the block-level scheduling strategy that we follow.
In addition to latency, we also compare the overall system throughput of Tally against the baseline GPU sharing systems. Tally demonstrates competitive performance relative to the baselines, attaining an average of $105.2\%$, $83.6\%$, $80.6\%$, and $80.3\%$ of the system throughput achieved by the four baseline systems, respectively. This shows the ability of Tally’s priority-aware scheduler to opportunistically schedule best-effort kernels while preserving latency requirements of high-priority tasks in shared execution environments.
Furthermore, we observe that MPS, MPS-Priority, and TGS exhibit substantial variation in performance degradation for high-priority inference tasks across different workload combinations. For instance, TGS experiences performance degradation ranging from $15.6\%$ to $751.7\%$. This variation in latency reveals the sensitivity of these systems to workload characteristics, such as the differences in kernel duration across various tasks. As these systems schedules at the kernel-level, their performance deteriorates more severely for workloads with longer kernels. In contrast, Tally’s block-level scheduling demonstrates robust performance against variations in kernel duration, effectively limiting the performance degradation of inference tasks to under $10\%$ across $70\%$ of the workload combinations (worst-case $23\%$ on ResNet50).
5.3. Traffic Load Sensitivity Analysis
We next evaluate the sensitivity of Tally’s performance isolation capabilities to varying traffic loads. For this experiment, we use BERT and Llama-2-7B as the inference tasks representative of low-latency ($3.93$ ms) and long-latency ($1.9$ s) workloads. Each of them is paired with three distinct training workloads: BERT, GPT2, and Whisper, chosen for their observed tendency to cause significant interference. Figure 5a (1) and (2) present a comparative analysis of the $99^{th}$-percentile latency for the inference tasks under Tally and TGS relative to the original latency as a function of GPU idle time (calculated as $100\%$ load) ranging from $10\%$ to $90\%$. The results demonstrate that both inference tasks under Tally exhibit latencies indistinguishable from their original performance consistently across varying loads. In contrast, TGS incurs latency slowdowns of up to 5.8×5.8$\times$5.8 × and 2.3×2.3$\times$2.3 × for BERT and Llama-2-7B, respectively. These findings demonstrate the robustness of Tally’s performance isolation capabilities for high-priority inference tasks, regardless of traffic load or the characteristics of the paired training workload.
We further analyze the relationship between Tally’s achievable throughput and traffic load. Figure 5a (3) and (4) illustrate the system throughput under Tally and TGS as a function of idle time. Both systems exhibit an upward trend in system throughput as the percentage of idle time increases, indicating their ability to leverage available resources for improved performance. During periods of high traffic load, there is a noticeable difference between the two systems. For instance, at $10\%$ idle time, the system throughput under Tally is $1.19$, whereas TGS achieves $1.65$. This difference is attributed to Tally’s strict priority enforcement policy, which executes best-effort tasks only when the high-priority task is inactive, sacrificing potential throughput gains for the preservation of the high-priority task’s latency. In contrast, TGS achieves higher throughput but at the cost of significant slowdowns in tail latency, as demonstrated earlier. Nevertheless, as idle time increases, the throughput difference between Tally and TGS diminishes. This convergence indicates that for scenarios where heavy underutilization is observed, Tally not only offers reliable performance isolation guarantees but also achieves competitive throughput compared to other GPU sharing systems.
Figure 6. (a) Scalability performance of Tally with respect to number of concurrent workloads. (b) Performance decomposition of Tally for BERT inference. (c) Latency and throughput under different turnaround latency threshold settings.
To further analyze the adaptive nature of Tally to varying traffic load, in Figure 5v, we plot the real-time traffic trace to a BERT inference task (a condensed version of the original MAF2 trace), and measure the real-time tail latency and throughput when co-located with another BERT training task. Figure 5v illustrates (1) the real-time traffic to a BERT inference task, (2) $99^{th}$ percentile latency of the inference server achieved by different GPU sharing systems and (3) the throughput of the co-located best-effort BERT training achieved by Tally compared to the original throughput of the training task. As shown in Figure 5b (2), the tail latency under Tally, represented by the purple line, closely aligns with the ideal latency depicted by the blue line, with negligible deviation throughout the simulation period. Whereas alternative approaches such as Time-Slicing, MPS, MPS-Priority, and TGS result in substantial slowdowns in latency as depicted. Additionally, Figure 5b (3) shows Tally’s ability to opportunistically adjust the throughput of the best-effort task in correspondence with the fluctuating traffic load to the high-priority inference task, as discussed in Section 4. As a result, Tally preserves over $68\%$ of the training task’s original throughput simultaneously running the high-priority inference workload on the same GPU without affecting its latency requirements.
5.4. Scalability with Number of Workloads
To demonstrate Tally’s ability to support multiple best-effort workloads while preserving the performance of the co-located high-priority job, we extend our experiments beyond the two-job scenario. We designed an experiment with a set of identical ResNet50 inference workloads, each operating at a low traffic load of $10\%$ and one of these workloads designated as the high-priority online inference task while the rest as best-effort offline inference tasks. We measure $99^{th}$-percentile latency of the high-priority task and the aggregated system throughput, measured in number of requests processed per minute, as we scale up the number of best-effort tasks. The results in Figure 6(a), reveal that the latency of the high-priority task remains consistent throughout the spectrum of scaling from $1$ to $10$ workloads. At the same time, the system throughput exhibits a steady increase as the number of workloads grows, until the GPU saturates at $8$ concurrent best-effort workloads. This shows that Tally can be particularly useful in scenarios where numerous jobs exhibit extremely low utilization, as Tally can efficiently pack these workloads together to maximize system throughput without compromising the performance of high-priority tasks.
5.5. Ablation study
Performance decomposition. We analyze the importance of Tally’s priority-aware scheduling algorithm and block-level transformation for performance isolation. Figure 6(b) compares the $99^{th}$-percentile latency of a high-priority BERT model when co-located with training workloads from the benchmark suite. Ideal denotes the original latency of the task executed in isolation. No-scheduling refers to the indiscriminate dispatching of kernels from both workloads upon arrival. Scheduling w/o transformation employs Tally’s priority-aware scheduling without kernel transformations, while Scheduling with transformation represents the full implementation of Tally, utilizing slicing and preemption mechanisms for block-level scheduling. The results reveal that No-scheduling incurs substantial latency overhead across all six job combinations, with the most severe case exhibiting a 30×30$\times$30 × slowdown when co-located with Whisper. This significant performance degradation is attributable to the absence of priority enforcement, resulting in considerable co-execution interference. The integration of Tally’s priority-aware kernel-level scheduling strategy somewhat mitigates this overhead. For ResNet50 and GPT2, this approach yields latencies close to ideal performance, with $8.0\%$ and $9.8\%$ slowdown respectively. However, for workloads such as Whisper and BERT, the latency slowdown can still reach up to 10×10$\times$10 ×. This difference in slowdown is attributed to the variation in kernel duration distribution. For instance, $99.3\%$ of kernels in ResNet50 complete in less than $0.1$ ms, whereas $5.6\%$ of kernels in Whisper exceed the duration of an entire BERT inference ($3.93$ ms). These prolonged kernels significantly delay the execution of high-priority kernels. Notably, with kernel transformation enabled, i.e., with full Tally, the latency overhead becomes negligible across all six training workloads, resulting in an average of $4.0\%$ slowdown ($6.2\%$ in the worst case). This shows the necessity of block-level scheduling in maintaining robust performance isolation guarantees.
Turnaround latency analysis. We examine the impact of the turnaround latency threshold (delineated in Section 4) by evaluating the tail latency of the BERT inference workload and the normalized throughput of the co-located training job on six models and across six distinct threshold values, ranging from $0.01$ ms to $10$ ms. The results, depicted in Figure 6 (c-i) and (c-ii), show that higher turnaround latency thresholds generally correspond to increased tail latency for the inference workload with only slight increase in the throughput. Based on these empirical results, we find that a turnaround latency threshold of $0.0316$ ms provides the best balance between latency and throughput, which is the default setting we use throughout experiments.
6. RELATED WORK
Many GPU sharing systems are designed with the goal of maximizing the throughput. Time-Slicing (GILMAN2021102234, ) and MPS (nvidia-mps, ) provided in NVIDIA GPUs enable transparent sharing of GPU resources in temporal and spatial manners. Besides vendor-provided options, many works (Ebird2019, ; gpupool2023, ; smgpu2016, ) have been built on the concept of co-executing kernels with complementary resource demands (e.g., compute vs. memory intensive) to further exploit resource parallelism in GPU SMs. For instance, Kernelet (zhong2013kernelet, ) and Slate (slate2019, ) perform such scheduling at runtime, while Kernel Fusion (fusion2010, ), HFUSE (li2022automatic, ), Tacker (tacker2022, ), and ISPA (ispa2023, ) employ source-to-source transformation to fuse kernels during compilation. Systems in this category fundamentally cannot support performance isolation as they prioritize the overall efficiency of the system over preservation of performance of certain high-priority workloads, thus falling short in production scenarios where tasks have SLAs.
Meanwhile, systems that are designed with performance isolation support, like Effisha (effisha2017, ) and GPES (gpes2015, ), utilize block-level preemption techniques to mitigate interference in concurrent execution but they require access to application source code. In contrast, Tally operates non-intrusively. AntMan (xiao2020, ) and TGS (tgs2023, ) leverage the iterative nature of DL training to adaptively limit the resources of low-priority tasks based on performance feedback from high-priority jobs. REEF (reef2022, ) implements a reset-based scheduling mechanism that achieves microsecond-scale kernel preemption latency on idempotent kernels. However, these works rely on specific task characteristics and cannot generalize to the wide variety of DL workloads. Conversely, Tally supports broad compatibility using task-agnostic scheduling and kernel transformations. Furthermore, Orion (orion2024, ) enhances performance isolation with kernel-level scheduling with consideration on each kernel’s compute and memory requirements. However, Orion requires offline profiling of workloads, which can be challenging to accomplish in cluster settings. Whereas Tally employs transparent profiling for priority-aware scheduling.
7. CONCLUSION
In this paper, we highlight the limitations of current GPU sharing systems and their inability to meet the versatile requirements of modern large-scale DL clusters. To address these limitations, we design Tally, a novel non-intrusive GPU sharing mechanism that offers robust performance isolation and universal applicability to diverse DL workloads. We empirically show that Tally can effectively ensure the performance standards of latency-critical tasks during shared execution, incurring a mere $7.2\%$ overhead on average on a diverse set of workload combinations, compared to 188.9% minimal overhead in state-of-the-art non-intrusive GPU sharing systems while offering competitive system throughput.